OpenStudy (anonymous):

Let n = 5. What is the value of 10 + 15 ÷ n – 2? A. 3 B. 6 C. 11 D. 15

OpenStudy (anonymous):

@flower2468

OpenStudy (flower2468):

alright so first you put the n in there so 10+15/5-2= next you do the divition first, 15/5=3 then you have 10+3-2 s next you do 10+3=13 then you have 13-2=...?

OpenStudy (anonymous):

11

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

My intenet is slow sorry

OpenStudy (anonymous):

next Let n = 2. What is the value of (32 + 6) ÷ n?

OpenStudy (flower2468):

SO you do 32+6 because of pharaences = 38/2=..?

OpenStudy (anonymous):

i will give medal if i get right

OpenStudy (anonymous):

38

OpenStudy (flower2468):

OH sorry, wrong you have to / it by 2 so 38/2 is 19

OpenStudy (anonymous):

19

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

Let q = 5. What is the value of q • (6 + 14)?

OpenStudy (flower2468):

alright so 6+14=20 and 20*5=..?

OpenStudy (anonymous):

100

OpenStudy (flower2468):

yes

OpenStudy (anonymous):

Thank you! I got a 100%

OpenStudy (flower2468):

you are welcome

OpenStudy (anonymous):

Thxs!

OpenStudy (anonymous):

I will text you if I need any more help

OpenStudy (anonymous):

Oh, It's more!

OpenStudy (flower2468):

okay

OpenStudy (anonymous):

Let x = 35. What is the value of x – 10? A. 25 B. 34 C. 35 D. 45

OpenStudy (flower2468):

okay so 35-10

OpenStudy (anonymous):

25

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

Let m = 26. What is the value of m + 11? A. 11 B. 15 C. 36 D. 37

OpenStudy (flower2468):

okay so 26+11

OpenStudy (anonymous):

37

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

Let s = 8. What is the value of 6 + 2s? A. 16 B. 22 C. 28 D. 64

OpenStudy (flower2468):

6+2*8 so 2*8=16 and 16+6=...?

OpenStudy (anonymous):

22

OpenStudy (flower2468):

corrrect

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let r = 3. What is the value of (10 + 17) ÷ r? A. 9 B. 17 C. 27 D. 30

OpenStudy (flower2468):

10+17=27/3=..?

OpenStudy (anonymous):

9

OpenStudy (flower2468):

corect

OpenStudy (anonymous):

Let b = 8. What is the value of 28 – b · 2? A. 10 B. 12 C. 24 D. 40

OpenStudy (anonymous):

Flower?

OpenStudy (flower2468):

yes sorry 8*2=16 and then 28-16..?

OpenStudy (anonymous):

12

OpenStudy (flower2468):

correect

OpenStudy (anonymous):

Let t = 5. What is the value of 6 · 4 + t · 3? A. 19 B. 39 C. 87 D. 133

OpenStudy (anonymous):

done

OpenStudy (flower2468):

5*3=15 and 4+15=...?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

hello?

OpenStudy (flower2468):

sorry yes

OpenStudy (anonymous):

Let t = 5. What is the value of 6 · 4 + t · 3? A. 19 B. 39 C. 87 D. 133

OpenStudy (anonymous):

I said 19

OpenStudy (flower2468):

okay so 6*4=24 and 5*3=15 so add those and get 24+15=..?

OpenStudy (anonymous):

39

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

Let s = 12. What is the value of 27 + 3s? Let s = 12. What is the value of 27 + 3s?

OpenStudy (flower2468):

so wow,okay, 3*12=36 and 36+27=...?

OpenStudy (anonymous):

63

OpenStudy (flower2468):

correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let h = 3. What is the value of (27 ÷ h) · 5?

OpenStudy (flower2468):

27/3=7 and 7*5=35

OpenStudy (anonymous):

35

OpenStudy (flower2468):

yes

OpenStudy (anonymous):

thank you! And sorry for asking so many questions!

OpenStudy (flower2468):

it okay, wha did you get

OpenStudy (anonymous):

100%

OpenStudy (flower2468):

great!:D

OpenStudy (anonymous):

(:

ILovePuppiesLol (ilovepuppieslol):

for future reference, please put this in the math section :)))

OpenStudy (khalid3166):

^^